OpenStudy (anonymous):

how do I Change 2x + 5y = –30 to slope-intercept form and State the slope and y-intercept??? someone plz help! I need somebody's help!

OpenStudy (anonymous):

You have to find y in terms of x for that..

OpenStudy (anonymous):

Firstly subtract 2x from both the sides and tell me what you get??

OpenStudy (anonymous):

2x + 5y = –30 5y = -2x - 30 y = (-2/5)x-6

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

Now compare your answer with the following equation: \(y = mx + c\) Tell me what is m there in your solution??

OpenStudy (anonymous):

Or simply tell me what is the coefficient of x there in your answer??

OpenStudy (anonymous):

5y = -2x +30

OpenStudy (anonymous):

@Nana3456 your first equation is right.. See, it is an example I am giving to you: See, if y = -6x + 3 Then here coefficient of x is -6 Similarly you have the equation: \[y = -\frac{2}{5}x - 6\] Tell me what is the coefficient of x here??

OpenStudy (anonymous):

:( I dont know ,,I always have problems with this kind of exercirses

OpenStudy (anonymous):

So Slope = \(\frac{-2}{5}\) and y - intercept = \(-6\)..

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

so that it?? :? :)

OpenStudy (anonymous):

And how do I state tthwe sloppe ans y intercept?

OpenStudy (anonymous):

Read the question carefully,, Simply state it as: Slope = \(\frac{-2}{5}\) and y - intercept = -6

OpenStudy (anonymous):

ok so I just leave it like that

OpenStudy (anonymous):

That is the answer you can't do more in that..

OpenStudy (anonymous):

ok thank oyou again:))

OpenStudy (anonymous):

Welcome dear.